Please show detailed work-up

Solve the IVP [j] = [-3 1] [:] + [6] [70] = [-]
Answer #1

As the method is not mentioned. I have solved it using laplace and partly without laplace.

Answer & Solve the IVP & [17:03:1811:47 ) :)] -6t zy - 3x + y we have X= -4x+ xlo)=4 y - y (0)=-6 Using the Laplace Transfo(54435+2) L(x) = -12-0 (5-0 +4 (5-1) 52 (s+i)(+2) L(), -12 -6 (S-1) +4 (5-1) S2 + LLY) = -12 8 - 6(S-1) 4 (3-1) (S+1)(3+2) s-3 y-y =-3x + from original equation & y-y = -3 (-8e-t- -Best+zae-2-15+3+) 117e-2t +45-9t 2 y-y - 24 e-teren 2 Linear +C st